# MDAgents: An Adaptive Collaboration of LLMs for Medical Decision-Making

### v2

@@ Abstract
Foundation models are becoming valuable tools in medicine. Yet despite their promise, the best way to leverage Large Language Models (LLMs) in complex medical tasks remains an open question. We introduce a novel multi-agent framework, named Medical Decision-making Agents (MDAgents) that helps to address this gap by automatically assigning a collaboration structure to a team of LLMs. The assigned solo or group collaboration structure is tailored to the medical task at hand, a simple emulation inspired by the way real-world medical decision-making processes are adapted to tasks of different complexities. We evaluate our framework and baseline methods using state-of-the-art LLMs across a suite of real-world medical knowledge and medical diagnosis benchmarks. MDAgents achieved the best performance in seven out of ten benchmarks on tasks requiring an understanding of medical knowledge and multi-modal reasoning, showing a significant improvement of up to 6.5% ($p$ < 0.05) compared to previous methods’ best performances. Ablation studies reveal that MDAgents effectively determines medical complexity to optimize for efficiency and accuracy across diverse medical tasks. Notably, the combination of moderator review and external medical knowledge in group collaboration resulted in an average accuracy improvement of 11.8%. Our code can be found at https://github.com/mitmedialab/MDAgents.
@@ 2 Related Works
LLMs have shown promise in a range of applications within the medical field [39, 59, 71, 91, 72, 14, 36, 86, 46]. They can answer questions from medical exams [41, 49], perform biomedical research [35], clinical risk prediction [36], and clinical diagnosis [63, 52]. Medical LLMs are also evaluated on generative tasks, including creating medical reports [75], describing medical images [77], constructing differentials [50], performing diagnostic dialogue with patients [73], and generating psychiatric evaluations of interviews [24]. To advance the capabilities of medical LLMs, two main approaches have been explored: (1) training with domain-specific data [28], and (2) applying inference-time techniques such as prompt engineering [63] and Retrieval Augmented Generation (RAG) [87]. While initial research has been concentrated on pre-training and fine-tuning with medical knowledge, the rise of large general-purpose LLMs has enabled training-free methods where models leverage their latent medical knowledge. For example, GPT-4 [56], with richer prompt crafting, surpasses the passing score on USMLE by over 20 points and with prompt tuning can outperform fine-tuned models including Med-PaLM [55, 54]. The promise of general-purpose models has thus inspired various techniques such as Medprompt and ensemble refinement to improve LLM reasoning [63], as well as RAG tools that use external resources to improve the factuality and completeness of LLM responses [87, 37]. Our approach leverages these techniques and the capabilities of general-purpose models while acknowledging that a solitary LLM [36, 46, 86] may not fully encapsulate the collaborative and multidisciplinary nature of real-world MDM. We thus emphasize joining multiple expert LLMs for effective collaboration in order to solve complicated medical tasks with greater accuracy.
@@ 3.1 Agent Roles
    High - Highly complex medical scenarios that demand extensive coordination and combined expertise from an ICT. These cases often involve multiple chronic conditions, complicated surgical or trauma cases, and complex decision-making that integrates various healthcare departments and multiple specialists.
@@ 4.2 Results
As depicted in Figure 4 and Table 2, MDAgents significantly outperforms (p < 0.05) both Solo and Group setting methods, showing best performance in 7 out of 10 medical benchmarks tested. This reveals the effectiveness of adaptive strategies integrated within our system, particularly when navigating through the text-only (e.g., DDXPlus where it outperformed the best performance of single-agent by 7.2% and multi-agent by 9.5%) and text-image datasets (e.g., Path-VQA, PMC-VQA and MIMIC-CXR). Our approach not only comprehends textual information with high precision but also adeptly synthesizes visual data, a pivotal capability in medical diagnostic evaluations.
Why Adaptive Decision-making Framework Works Well.
Figure 3: Experiment with the MedQA dataset (N=25 randomly sampled questions). (a) LLM’s capability to classify complexity. (b-d) Evaluating 25 medical problems by solving each problem 10 times at each difficulty level.
@@
We hypothesize that an LLM, functioning as a classifier or a generalist doctor, will select the optimal complexity level for each MDM problem. This hypothesis is supported by Figure 3, which illustrates that the model appropriately matches the complexity levels; low, moderate, and high of the given problem. To determine this, we assessed the accuracy of solutions across various difficulty levels. Specifically, we evaluated 25 medical problems by solving each problem 10 times at each difficulty level. By measuring the success rate, we aimed to identify the difficulty level that yielded the highest accuracy. This rigorous approach ensures that the LLM’s classification of problem difficulty aligns with the most effective and accurate medical solutions, thereby optimizing the application of medical expertise to each question.
@@ D.4 Complexity Assignment and Collaborative Settings
Gemini-Pro(Vision) Majority Voting 52.3 $\pm$ 15.3 73.5 $\pm$ 4.9 47.0 $\pm$ 1.4 44.0 $\pm$ 4.2 47.9 $\pm$ 6.6
Weighted Voting 54.3 $\pm$ 16.9 64.6 $\pm$ 6.5 42.0 $\pm$ 3.2 43.6 $\pm$ 3.9 43.2 $\pm$ 2.0
Borda Count 67.0 $\pm$ 27.7 77.3 $\pm$ 9.3 37.0 $\pm$ 1.8 46.1 $\pm$ 3.2 44.6 $\pm$ 4.5
MedAgents [68] 43.0 $\pm$ 2.7 80.5 $\pm$ 1.9 40.8 $\pm$ 2.9 50.5 $\pm$ 3.6 38.7 $\pm$ 1.5
Multi-agent Reconcile [10] 68.4 $\pm$ 7.4 90.6 $\pm$ 2.5 60.7 $\pm$ 5.7 59.5 $\pm$ 8.7 33.3 $\pm$ 3.4
DyLAN [48] 56.4 $\pm$ 2.9 75.3 $\pm$ 4.6 60.1 $\pm$ 3.1 57.3 $\pm$ 6.1 38.7 $\pm$ 1.2
* T : text-only, I : image+text, V : video+text
* All experiments were tested with 3 random seeds
Table 11: Accuracy (%) on Medical benchmarks with Our ([Uncaptioned image]) method. Bold represents the best performance for each benchmark and model.
Method Medical Knowledge Retrieval Datasets
MedQA T PubMedQA T Path-VQA I T PMC-VQA I T MedVidQA V T
GPT-3.5 64.0 $\pm$ 1.6 66.0 $\pm$ 5.7 - - -
GPT-4(V) 88.7 $\pm$ 4.0 75.0 $\pm$ 1.0 65.3 $\pm$ 3.9 56.4 $\pm$ 4.5 -
Gemini-Pro(Vision) 57.4 $\pm$ 1.8 71.0 $\pm$ 1.6 72.0 $\pm$ 2.3 62.2 $\pm$ 7.6 56.2 $\pm$ 6.7
Method Clinical Reasoning and Diagnostic Datasets
DDXPlus T SymCat T JAMA T MedBullets T MIMIC-CXR I T
GPT-3.5 62.5 $\pm$ 6.7 85.7 $\pm$ 3.0 48.6 $\pm$ 6.5 55.3 $\pm$ 4.3 -
GPT-4(V) 77.9 $\pm$ 2.1 93.1 $\pm$ 1.0 70.9 $\pm$ 0.3 80.8 $\pm$ 1.7 55.9 $\pm$ 9.1
Gemini-Pro(Vision) 59.2 $\pm$ 1.2 65.0 $\pm$ 6.1 47.0 $\pm$ 2.2 42.9 $\pm$ 3.4 48.1 $\pm$ 5.8
* T : text-only, I : image+text, V : video+text
* All experiments were tested with 3 random seeds
Table 12: Accuracy (%) on Medical benchmarks with Solo/Group/Adaptive settings with increased number of samples (N=100). All benchmarks except for MedVidQA (Gemini 1.5 Flash) were evaluated with GPT-4o mini.
Category Method MedQA T PubMedQA T Path-VQA I T PMC-VQA I T MedVidQA V T
Single-agent Zero-shot 75.0 54.0 58.0 48.0 50.0
Few-shot 77.0 55.0 58.0 50.0 51.0
+ CoT 78.0 50.0 59.0 52.0 53.0
+ CoT-SC 79.0 51.0 60.0 53.0 53.0
ER 76.0 51.0 61.0 51.0 52.0
Medprompt 79.0 58.0 60.0 54.0 53.0
Multi-agent (Single-model) Majority Voting 79.0 68.0 63.0 52.0 54.0
Weighted Voting 80.0 68.0 64.0 51.0 55.0
Borda Count 81.0 69.0 62.0 50.0 52.0
MedAgents 80.0 69.0 55.0 52.0 50.0
Meta-Prompting 82.0 69.0 56.0 49.0 -
[…]
AutoGen 65.0 63.0 45.0 40.0 -
DyLAN 68.0 67.0 42.0 48.0 -
Adaptive MDAgents (Ours) 87.0 71.0 60.0 55.0 56.0
Category Method DDXPlus T SymCat T JAMA T MedBullets T MIMIC-CXR I T
Single-agent Zero-shot 53.0 84.0 57.0 49.0 38.0
Few-shot 60.0 87.0 58.0 52.0 33.0
+ CoT 66.0 84.0 55.0 64.0 33.0
+ CoT-SC 68.0 84.0 57.0 60.0 40.0
ER 76.0 80.0 56.0 59.0 43.0
Medprompt 70.0 84.0 62.0 60.0 43.0
Multi-agent (Single-model) Majority Voting 53.0 82.0 56.0 59.0 54.0
Weighted Voting 52.0 86.0 56.0 56.0 52.0
Borda Count 53.0 86.0 56.0 59.0 51.0
MedAgents 56.0 80.9 51.0 58.0 40.9
Meta-Prompting 53.0 79.0 56.0 51.0 48.0
[…]
AutoGen 47.0 87.0 53.0 55.0 47.0
DyLAN 54.0 84.0 55.0 57.0 42.0
Adaptive MDAgents (Ours) 75.0 89.0 59.0 67.0 56.0
* T : text-only, I : image+text, V : video+text
Figure 8: Complexity Distribution for each dataset classified by GPT-4(V) and Gemini-Pro (Vision) (for MedVidQA). The plot illustrates the varying levels of medical complexity across datasets, reflecting the diverse nature of medical question answering, diagnostic reasoning, and medical visual interpretation tasks. For instance, MedQA is categorized under Medical Knowledge Retrieval due to their focus on text-based questions and literature synthesis, while MIMIC-CXR, categorized under Clinical Reasoning and Diagnostic tasks, shows a high complexity distribution due to the need for interpreting detailed radiographic images (See Section in Section 4.1 for the task categorization)
Figure 9: Top-3 most recruited medical experts in each benchmark. The alignment between the dataset characteristics and the recruited experts is evident in several cases. For instance, MIMIC-CXR, which features chest x-ray images, predominantly recruits Radiologists, Pulmonologists, and experts in Biomedical Informatics due to their expertise in interpreting medical imaging.
Figure 10: Simplified agent structure examples assigned during the expert recruitment process ranging from (a) A Primary Care Clinician (PCC), (b) Multi-disciplinary Team (MDT), (C) MDT w/ hierarchy to (d) Integrated Care Team (ICT).
Algorithm 1 Adaptive Medical Decision-making Framework
1:Problem $Q$
2:$Complexity\leftarrow\textsc{ComplexityCheck}(Q)$triangleright Determine the complexity of the medical query
3:if $Complexity=low$ then
4:    $Agent\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit a Primary Care Clinician agent
5:    $ans\leftarrow Agent(Q)$
6:else if $Complexity=moderate$ then
7:    $MDT\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit a Multi-disciplinary Team
8:    $Agent\leftarrow\textsc{Recruit}(Q,Complexity,MDT)$
9:    $r\leftarrow 0$
10:    $Consensus\leftarrow\text{False}$
11:    $Interaction\leftarrow[]$
12:    while $r\leq R$ and not $Consensus$ do
13:        $Consensus,Log\leftarrow\textsc{CollaborativeDiscussion}(Q,MDT)$triangleright Iterative discussions
14:        if not $Consensus$ then
15:           for all $Agent\in MDT$ do
16:               $Feedback\leftarrow Moderator(Interaction,Agent)$triangleright Review and provide feedback
17:               $Agent.\textsc{Update}(Feedback)$triangleright Update the feedback
18:           end for
19:           $Interaction\leftarrow Interaction+[Log]+[Feedback]$
20:        end if
21:        $r\leftarrow r+1$
22:    end while
23:    $ans\leftarrow Agent(Q,Interaction)$triangleright Moderator agent makes the final decision
24:else
25:    $ICT\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit an Integrated Care Team
26:    $Reports\leftarrow[]$
27:    for $Team\in ICT$  do
28:        $Report\leftarrow\textsc{GenerateReport}(Q,Team)$triangleright Each Team curates a report
29:        $Reports\leftarrow Reports+[Report]$
30:    end for
31:    $ans\leftarrow Agent(Q,Reports)$triangleright Final decision made
32:end if
33:return $ans$
Appendix E Case Study
E.1 Medical Decision Making Case Studies
MDM requires efforts of both individual expertise and collaboration to navigate the complexities of patient care. Clinicians often face challenging scenarios that necessitate a comprehensive approach, integrating insights from various specialties to arrive at the best possible outcomes.
E.1.1 Real-World Medical Cases
Below are the real-world example cases that could be classified as low, moderate, to high complexity cases.
Case 1: Adjusting Medication Dosage for Chronic Disease (Low Complexity)
A 55-year-old female patient with type 2 diabetes visits her PCP for a routine check-up. The patient has been taking 500 mg of metformin orally twice a day and has been adhering to a low-carbohydrate diet. Upon testing with fasting glucose level, the glucose level is above normal. PCP reviewed the current medical dosage and increased the dosage to manage the blood glucose level of the patient.
Case 2: Differential Diagnosis in the Emergency Department (Moderate Complexity)
A 40-year-old male patient arrives at the emergency department (ED) with a high fever, severe headache, and muscle pain, raising concerns about a potential infectious disease. The ED physician conducts an initial examination but recognizes the need for a more detailed evaluation to identify the underlying cause. The patient is referred to the infectious disease department for further assessment. An infectious disease specialist, along with the ED physician, reviews the patient’s symptoms, travel history, and recent exposures. They collaborate on ordering specific diagnostic tests, including blood cultures and imaging studies. Through this teamed decision-making process, they diagnose the patient with dengue fever and promptly initiate appropriate antiviral treatment.
Case 3: Managing Adverse Responses to Medication in Chronic Disease (High Complexity)
A 60-year-old female patient with chronic heart failure has been experiencing new symptoms of shortness of breath and mild fever, suggesting either a complication due to her chronic heart failure or a new infection. The urgent care doctor identifies the severity of the situation and promptly refers the patient to the emergency department of a large hospital, where the patient has triaged to see a cardiologist and an infectious disease doctor for specialized care. The team conducts a detailed review of the patient’s medication history and current symptoms, does a physical exam to listen to lung sounds, and orders a few exams including labs, a chest x-ray, echocardiogram, and electrocardiogram. The team identifies that the patient has pulmonary effusion and upper respiratory viral infection.
E.1.2 Medical Cases from MedQA Dataset
Now, let us look at the cases from the MedQA [34] dataset that illustrate either individual PCP or teamed decision-making is crucial in managing medical conditions, ranging from low to high complexity levels of potential cases. These examples highlight the importance of checking the complexity of the case for proper management.
Case 4: Diagnosis by PCP
The case below with the “Low Complexity” header is classified as low complexity by a medical doctor. In this case, a PCP can answer this question without consulting a gastroenterologist. The diagnosis of gastric cancer and management based on the manifestation of the disease, that has been described in this question and beyond should be from a gastroenterologist. However, PCPs are expected to have the basic scientific and pathophysiological knowledge that is related to gastric cancer and use that knowledge to solve this problem.
Question: A 70-year-old man comes to the physician because of a 4-month history of epigastric pain, nausea, and weakness. He has smoked one pack of cigarettes daily for 50 years and drinks one alcoholic beverage daily. He appears emaciated. He is 175 cm (5 ft 9 in) tall and weighs 47 kg (103 lb); BMI is 15 kg/m². He is diagnosed with gastric cancer. Which of the following cytokines is the most likely direct cause of this patient’s examination findings? Answer: A) TGF-$\beta$ B) IL-6 C) IL-2 D) TNF-$\beta$ Low Complexity
Case 5: Diagnosis and management by single Pediatric Endocrinologist
The case below with the “Moderate Complexity” header is classified as moderate complexity by the medical doctor. In this case, a pediatric endocrinologist (specialist) alone can diagnose a patient and have a treatment plan. Note that this patient could have been referred to this pediatric endocrinologist by a PCP who is regularly seeing this patient.
Question: A 5-year-old girl is brought to the clinic by her mother for excessive hair growth. Her mother reports that for the past 2 months she has noticed hair at the axillary and pubic areas. She denies any family history of precocious puberty and reports that her daughter has been relatively healthy with an uncomplicated birth history. She denies any recent illnesses, weight change, fever, vaginal bleeding, pain, or medication use. Physical examination demonstrates Tanner stage 4 development. A pelvic ultrasound shows an ovarian mass. Laboratory studies demonstrate an elevated level of estrogen. What is the most likely diagnosis? Answer: A) Granulosa cell tumor B) Idiopathic precocious puberty C) McCune-Albright syndrome D) Sertoli-Leydig tumor Moderate Complexity
Case 6: Diagnosis and management by multidisciplinary team
The case below is of a “High Complexity” patient primarily having neurological symptoms but with problems with vision, which requires a neurologist to consult to ophthalmology department for further evaluation.
Question: A 63-year-old woman presents to her primary-care doctor for a 2-month history of vision changes, specifically citing the gradual onset of double vision. Her double vision is present all the time and does not get better or worse throughout the day. She has also noticed that she has a hard time keeping her right eye open, and her right eyelid looks ’droopy’ in the mirror. Physical exam findings during primary gaze are shown in the photo. Her right pupil is 6 mm and poorly reactive to light. The rest of her neurologic exam is unremarkable. Laboratory studies show an Hb A1c of 5.0%. Which of the following is the next best test for this patient? Answer: A) Direct fundoscopy B) Intraocular pressures C) MR angiography of the head D) Temporal artery biopsy High Complexity
E.2 Cases Studies with MDAgents
In this section, we provide two examples of our framework with moderate (Figure 11) and high (Figure 12) complexity in PMC-VQA (image+text) and DDXPlus (text-only) respectively. These case studies reveals how our framework provides an environment for agents to collaborate, gather information, moderate and make final decisions in complex medical scenarios.
Figure 11: Illustration of our proposed framework in moderate complexity setting. Given a medical query (image + text) the framework performs reasoning in five steps: (i) complexity check, (ii) expert recruitment, (iii) initial assessment, (iv) collaborative discussion, and (v) review and final decision-making. Green text represents the correct answer and the Red text represents the incorrect answer.
Figure 12: Illustration of our proposed framework in high complexity setting. Given a medical query (text-only) the framework performs reasoning in four steps: (i) complexity check, (ii) expert recruitment, (iii) information gather and report generation, (iv) final decision. Green text represents the correct answer.

### v3

@@ Abstract
Foundation models are becoming valuable tools in medicine. Yet despite their promise, the best way to leverage Large Language Models (LLMs) in complex medical tasks remains an open question. We introduce a novel multi-agent framework, named Medical Decision-making Agents (MDAgents) that helps to address this gap by automatically assigning a collaboration structure to a team of LLMs. The assigned solo or group collaboration structure is tailored to the medical task at hand, a simple emulation inspired by the way real-world medical decision-making processes are adapted to tasks of different complexities. We evaluate our framework and baseline methods using state-of-the-art LLMs across a suite of real-world medical knowledge and medical diagnosis benchmarks, including a comparison of LLMs’ medical complexity classification against human physicians111 Appendix F contains a detailed comparison results between human physicians and LLMs.. MDAgents achieved the best performance in seven out of ten benchmarks on tasks requiring an understanding of medical knowledge and multi-modal reasoning, showing a significant improvement of up to 4.2% ($p$ < 0.05) compared to previous methods’ best performances. Ablation studies reveal that MDAgents effectively determines medical complexity to optimize for efficiency and accuracy across diverse medical tasks. Notably, the combination of moderator review and external medical knowledge in group collaboration resulted in an average accuracy improvement of 11.8%. Our code can be found at https://github.com/mitmedialab/MDAgents.
@@ 2 Related Works
LLMs have shown promise in a range of applications within the medical field [40, 63, 75, 96, 76, 14, 37, 90, 48]. They can answer questions from medical exams [43, 52], perform biomedical research [36], clinical risk prediction [37], and clinical diagnosis [67, 55]. Medical LLMs are also evaluated on generative tasks, including creating medical reports [79], describing medical images [81], constructing differentials [53], performing diagnostic dialogue with patients [77], and generating psychiatric evaluations of interviews [24]. To advance the capabilities of medical LLMs, two main approaches have been explored: (1) training with domain-specific data [28], and (2) applying inference-time techniques such as prompt engineering [67] and Retrieval Augmented Generation (RAG) [92]. While initial research has been concentrated on pre-training and fine-tuning with medical knowledge, the rise of large general-purpose LLMs has enabled training-free methods where models leverage their latent medical knowledge. For example, GPT-4 [60], with richer prompt crafting, surpasses the passing score on USMLE by over 20 points and with prompt tuning can outperform fine-tuned models including Med-PaLM [59, 58]. The promise of general-purpose models has thus inspired various techniques such as Medprompt and ensemble refinement to improve LLM reasoning [67], as well as RAG tools that use external resources to improve the factuality and completeness of LLM responses [92, 38]. Frameworks like MEDIQ [49] and UoT [33] advance LLM reliability in clinical settings by enhancing information-seeking through adaptive question-asking and uncertainty reduction, supporting more realistic diagnostic processes. Our approach leverages these techniques and the capabilities of general-purpose models while acknowledging that a solitary LLM [37, 48, 90] may not fully encapsulate the collaborative and multidisciplinary nature of real-world MDM. We thus emphasize joinging multiple expert LLMs for effective collaboration in order to solve complicated medical tasks with greater accuracy.
@@ 3.1 Agent Roles
    High - Complex medical scenarios that demand extensive coordination and combined expertise from an ICT. These cases often involve multiple chronic conditions, complicated surgical or trauma cases, and decision-makings that integrates specialists from different healthcare departments.
@@ 4.2 Results
As depicted in Figure 4 and Table 2, MDAgents significantly outperforms (p < 0.05) both Solo and Group setting methods, showing best performance in 7 out of 10 medical benchmarks tested. This reveals the effectiveness of adaptive strategies integrated within our system, particularly when navigating through the text-only (e.g., DDXPlus where it outperformed the best performance of single-agent by 5.2% and multi-agent by 9.5%) and text-image datasets (e.g., Path-VQA, PMC-VQA and MIMIC-CXR). Our approach not only comprehends textual information with high precision but also adeptly synthesizes visual data, a pivotal capability in medical diagnostic evaluations.
Why Do Adaptive Decision-making Framework Work Well?
Figure 3: Experiment with the MedQA dataset (N=25 randomly sampled questions). (a) LLM’s capability to classify complexity. (b-d) Evaluating 25 medical problems by solving each one 10 times at various complexity levels. The x-axis represents the accuracy achieved for each problem, while the y-axis shows the number of problems that reached that level of accuracy.
@@
We hypothesize that an LLM, functioning as a classifier, will select the optimal complexity level for each MDM problem. This hypothesis is supported by Figure 3, which illustrates that the model appropriately matches the complexity levels; low, moderate, and high of the given problem. To determine this, we assessed the accuracy of solutions across various difficulty levels. Specifically, we evaluated 25 medical problems by repeating each problem for 10 times at each difficulty level. By measuring the success rate, we aimed to identify the difficulty level that yielded the highest accuracy. This approach ensures that the LLM’s complexity classification aligns with the most effective and accurate medical solutions, thereby optimizing the application of medical expertise to each question.
@@ D.4 Complexity Assignment and Collaborative Settings
Weighted Voting